# MelCap: A Unified Single-Codebook Neural Codec for High-Fidelity Audio Compression

### v1

@@ 3 Methods
GAN-based end-to-end codecs require dedicated discriminators and multiple codenooks to improve the waveform details, high-frequency components, and phase synchronization, but training these discriminators is time-consuming and convergence is often slow Wu et al. (2023). Consequently, we propose a two-stage codec, where the first stage focuses on mel-spectrogram reconstruction with metric losses, and the second stage incorporates a discriminator to recover high-fidelity waveform from mel discrete tokens. This architecture significantly improves training efficiency, enabling our second-stage model to converge within only 50 epochs.
@@ 4 Results
DataSets. The first-version model is trained on the AudioSet dataset, using the entire training subset (bal train). The AudioSet covers a wide range of sounds, including human and animal vocalizations, musical instruments and genres, as well as common everyday environmental noises. We train a second version adding speech dataset hq-conversations. All audio files are kept at their original sampling rate of 44 kHz. For each audio sample, mel-scaled spectrograms are computed with the following parameters: FFT size $n_{fft}$=1024, hop size $hop_{n}$ =256, and 96 Mel bins.
@@
Baseline Methods. Our proposed model is compared against DAC  Kumar et al. (2023), SNAC Siuzdak et al. (2024), Mel Codec  Langman et al. (2025), NVIDIA NeMo Audio Codec and WavTokenizer  Ji et al. (2025). For all baselines, we use the officially released pretrained checkpoints— the 24kHz version for wavTokenizer and 44 kHz versions for other methods, which are publicly available online.
@@ 4.1 Evaluation
We evaluate our models using four primary objective metrics, VISQOL, LSD, Mel Distance, STFT Distance, and two additional reference metrics, UTMOS and V/UV F1. The primary metrics assess spectral and perceptual fidelity, while the reference metrics are included for informational purposes, as they are designed for speech and may be less reliable in general audio tests.

### v2

@@ 3 Methods
GAN-based end-to-end codecs require dedicated discriminators and multiple codebooks to improve the waveform details, high-frequency components, and phase synchronization, but training these discriminators is time-consuming and convergence is often slow Wu et al. (2023). Consequently, we propose a two-stage codec, where the first stage focuses on mel-spectrogram reconstruction with metric losses, and the second stage incorporates a discriminator to recover high-fidelity waveform from mel discrete tokens. This architecture significantly improves training efficiency, enabling our second-stage model to converge within only 50 epochs.
@@ 4 Experiments
DataSets. The model is trained on the AudioSet dataset, using the entire balanced training subset (bal train), along with the HQ-Conversations dataset Magic Data (2024). The AudioSet covers a wide range of sounds, including human and animal vocalizations, musical instruments and genres, as well as common everyday environmental noises. We evaluate the in-domain reconstruction performance of the codec on the AudioSet test set. We evaluate the out-of-domain reconstruction performance of the codec on a subset of the music genre dataset  Lewtun (2022), which consists of 1,269 randomly selected 10-second clips. All audio files are kept at their original sampling rate of 44 kHz. For each audio sample, mel-scaled spectrograms are computed with the following parameters: FFT size $n_{fft}$=1024, hop size $hop_{n}$ =256, and 96 Mel bins.
@@
Baseline Methods. Our proposed model is compared against DAC  Kumar et al. (2023), SNAC Siuzdak et al. (2024), Spectral Codec  Langman et al. (2025), NVIDIA NeMo Audio Codec and WavTokenizer  Ji et al. (2025). For all baselines, we use the officially released pretrained checkpoints— the 24kHz version for WavTokenizer and 44 kHz versions for other methods, which are publicly available online. For a fair comparison, we additionally trained a DAC model with a smaller number of quantizers (four codebooks) to match the bitrate of our proposed method.
@@ 4.1 Evaluation
We evaluate our models using four primary objective metrics, VISQOL, LSD, Mel Distance, STFT Distance, and two additional reference metrics, UTMOS and V/UV F1. We also performed subjective evaluation with MUSHRA. The primary metrics assess spectral and perceptual fidelity, while the reference metrics are included for informational purposes, as they are designed for speech and may be less reliable in general audio tests.